An object slides down an inclined plane from an initial height of 30 meters.Its velocity at the foot of the plane is 20m/s.
The percentage of its initial potential energy which is deposited as heat is?
1
Expert's answer
2020-04-27T10:00:19-0400

The initial potential energy

"PE_i=mgh"

The final kinetic energy

"KE_f =\frac{mv^2}{2}"

The heat losses

"Q=PE_i-KE_f=mgh\left(1-\frac{v^2}{2gh}\right)"

Hence, the percentage of its initial potential energy which is deposited as heat

"\frac{Q}{PE_i}=1-\frac{v^2}{2gh}=1-\frac{20^2}{2\times 9.8\times 30}=0.32=32\%"